Given:

Titrant: HCl

Analyte: NH3

Find-> pH at HALF-WAY point

Find--> Volume at HALF-WAY point

Find--> pH at EQUIVALENCE POINT

Find--> Volume at EQUIVALENCE POINT


Data

Titration Lab

HCl + NH3

Volume of 0.10 M NH3 added (mL)--> 8.05 mL

interval=0.5 mL

Volume 0.10 M HCl Added (mL)

0.00

0.50

1.00

1.50

2.00

2.50

3.00

3.50

4.00

4.50

5.00

5.75

6.00

6.50

7.00

7.50

8.00

9.00

PH

10.05

10.16

10.00

9.87

9.76

9.61

9.53

9.40

9.31

9.20

9.10

8.91

8.74

8.30

7.60

6.79

3.55

3.18



1
Expert's answer
2021-04-19T05:14:04-0400

pH = - log [H3O+].


pH at HALF-WAY = pKab


pOH = -log 2.79x10-6 = 5.55


pH = 14 - 5.55 = 8.45